Solve each system by the substitution method:
x = -3y + 2
7y = 3x

Answer by Alan3354(69443)

Solve each system by the substitution method:
x = -3y + 2
7y = 3x
-------------
x is already given in terms of y in the 1st eqn, so sub into the 2nd.
7y = 3*(-3y+2)
7y = -9y+6
16y = 6
y = 3/8
--------------
x = -3y+2
x = (-9/8)+2
x = 7/8